how much metal that is 75% iron must be mixed with metal that is 60%iron in order to get 200 kilograms of a metal that is 69%iron
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

how much metal that is 75% iron must be mixed with metal that is 60%iron in order to get 200 kilograms of a metal that is 69%iron
------
Equations:
x + y = 200 kg
0.75x + 0.60y = 0.69*200
---------------------------------
Modify for elimination::
75x + 75y = 75*200
75x + 60y = 69*200
-----------------------
Subtract and solve for "y"::
15y = 6*200
y = (2/5)200
y = 2*40 = 80 kg (amt. of 75% iron needed)
x = 200-y = 120 kg (amt. of 60% iron needed)
----------------
Cheers,
Stan H.
-----------

Answer by addingup(3677)

amount of metal: 200, concentration: 69%
amount at 75% iron: x
amount at 60% iron: 200 - x
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
0.75x + 0.60(200 - x) = 0.69(200)
0.75x + 120 - 0.60x = 138
0.15x = 18
x = 120
You need 120 kg of 75% and 200-120 = 80 kg of 60%